A firm’s balance sheets for year-end 2016 and 2017 contain the following data. All items are in millions of dollars.

Dec. 31, 2016 Dec. 31, 2017
Accounts receivable 33.6 36.8
Inventories 26.6 30.8
Accounts payable 13.6 26.8

What happened to investment in net working capital during 2017? (Enter your answer in millions.)
Answer #1

Increase in Accounts receivable =36.8-33.6 =3.2

Increase in inventories =30.8-26.6 =4.2

Increase in accounts payable =26.8-13.6 = 13.2

Net working capital inflow=3.2+4.2-13.2 =5.8 million